A firm wants to minimize cost when Q=160=2√KL ;r=4 k=2

a) write out the Lagrangian function

b) Find K* and L*

c).Is the second-order sufficient condition for minimum satisfied?


1
Expert's answer
2021-10-02T10:49:29-0400

a) The Lagrangian can be written as

"l=4K+2L+\lambda[80-\sqrt{KL}]"

where λ is the Lagrange multiplier.


b)

The first order conditions can be written as

"\frac{\delta l}{\delta K}=4-\lambda \frac{ 1}{2}K^{\frac{-1}{2}}L^{\frac{1}{2}}=0\\or\\8=\lambda \sqrt{\frac{L}{K}}.....(1)\\

\frac{\delta l}{\delta K}=2-\lambda \frac{ 1}{2}K^{\frac{1}{2}}L^{\frac{-1}{2}}=0\\
or\\
4=\lambda \sqrt{\frac{K}{L}}.......(2)\\

\frac{\delta l}{\delta K}=80-\sqrt{KL}.....(3)"


Now, (1)/(2) gives -

2 = L/K or L = 2K

 

Substituting this value in (3) gives -

"2 =\frac{ L}{K}\\ or\\ L = 2K"

Substituting this value in (3) gives -


"\sqrt{2K^2}=80\\or\\
K^*=\frac{80}{\sqrt{2}}=56.5\\
L^*=\frac{160}{\sqrt{2}}=113.14"


c.

Second order conditions are 

"\frac{\delta ^2 l}{\delta K^2}=-\frac{\lambda}{2}(-\frac{1}{2})K^{\frac{-3}{2}}L^{\frac{1}{2}}\\=\frac{\lambda \sqrt{L}}{4K^{\frac{-3}{2}}}>0"

"\frac{\delta ^2 l}{\delta L^2}=-\frac{\lambda}{2}(-\frac{1}{2})K^{\frac{1}{2}}L^{\frac{-3}{2}}\\

=\frac{\lambda \sqrt{K}}{4L^{\frac{3}{2}}}>0"

Since second order partial derivatives are positive, hence we can conclude that condition for cost minimisation is satisfied.